# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ II Waveforms for eccentric equatorial inspirals
where $\mu=m_{1}m_{2}/M$ is the reduced mass of the binary, $M=m_{1}+m_{2}$ is its total mass, $d_{\mathrm{L}}$ is the luminosity distance to the source, and $(\theta,\phi)$ are, respectively the source-frame polar and azimuthal viewing angles of the observer. Our notation for the masses in our waveform model deviates from the majority of previous adiabatic waveform models in the GSF literature, e.g., Fujita:2020zxe ; Hughes:2021exa ; Isoyama:2021jjd ; Katz:2021yft ; Nasipak:2023kuf ; Khalvati:2024tzz , which define $\mu=m_{2}$ and $M=m_{1}$. Instead, our mass convention is chosen to align with what is commonly used when modelling comparable-mass binaries, e.g., Blanchet:2013haa ; Blanchet:2023bwj ; Pompili:2023tna ; Albanesi:2025txj . Further discussion of this change in convention is provided in Appendix A.
In the following section, we review how Eq. 1 is constructed in the few framework, highlighting where assumptions made for non-spinning systems must be modified in the extension of this framework to spinning systems.
@@ II.1 Inspiral trajectories into spinning black holes
In order to compute the amplitude and phase of each mode in Eq. 1, we must first obtain the orbital trajectory of the CO by integrating its equations of motion. Operating in the multiscale framework at adiabatic order Hinderer:2008dm ; Hughes:2016xwf ; pound2022black ; Isoyama:2021jjd , we parameterize the inspiral trajectory at each moment in time by three quasi-Keplerian orbital elements — $p(t)$ (semi-latus rectum), $e(t)$ (eccentricity) and $x_{I}(t)$ (cosine of the orbital inclination)— along with the fundamental orbital phases $\Phi_{A}(t)$ with $A=\{r,\theta,\phi\}$ (up to initial conditions). In this work, we specifically choose these elements such that they satisfy the geodesic relations between semi-latus rectum, eccentricity, inclination and the (dimensionless) orbital frequencies $\hat{\Omega}_{A}(a,p,e,x_{I})$, as described in Lynch:2024hco (built on works Glampedakis:2002ya ; Schmidt:2002qk ; Fujita:2009bp ; KerrGeodesicsPackage ), so that it adheres to the convention set out in the LISA convention document lisa_rosetta_stone_2025 . [^4] Some works treat $p$ as dimensionless (e.g., Fujita:2009bp ), while others take $p$ to scale with $m_{1}$ (e.g., Lynch:2024hco ). In this work, we make all quantities, including $p$, dimensionless and explicitly introduce the dependence on mass through the equations of motion. We discuss our choice of mass convention in Appendix A. The equations of motion then describe how these quantities evolve with respect to time. At adiabatic order, the evolution of the orbital elements is driven solely by the orbit-averaged back-reaction due to GW emission, and the evolution of the fundamental phases is equivalent to that of the tangent geodesic (which changes with respect to time due to the slowly evolving orbital elements). The adiabatic equations of motion are therefore a system of six ordinary differential equations (ODEs) that must be solved numerically. Three of these ODEs are associated with the evolution of the (dimensionless) orbital elements $\alpha\in\{p,e,x_{I}\}$:
@@
These phases can then be related to the $\{r,\theta,\phi\}$ Boyer-Lindquist coordinates of the secondary using the relations found in Ref. Lynch:2024hco . The phase evolution of each mode in Eq. 1 is then
@@
However, as only modes with $k=0$ contribute to the waveform for equatorial inspirals, we may write $\Phi_{mn}=m\Phi_{\phi}+n\Phi_{r}$ and ignore $\Phi_{\theta}$. By combining Eqs. 2 and 3, and integrating over a radiation reaction timescale ($\sim M/\nu$), we can see that the truncation in the evolution of the orbital elements leads to an error in the orbital phase of $O(\nu^{0})$ Hinderer:2008dm  [^5] In this work, we are ignoring the GSF effects of transient resonances Flanagan:2010cd , which first enter the inspiral evolution at $\mathcal{O}(\nu^{1/2})$ past the adiabatic order vandeMeent:2013sza ; Berry:2016bit ; Lynch:2024ohd . Transient resonances for equatorial eccentric inspirals are even weaker, entering at $\mathcal{O}(\nu^{1})$ (i.e. first post-adiabatic order) vandeMeent:2014raa and can therefore be neglected in our adiabatic model.. This is sufficient to demonstrate the few framework for IMRIs and EMRIs, and to obtain the quantitive results of the studies performed in this work. Once post-adiabatic corrections to the evolution of the orbital elements are known, they can easily be incorporated as additional forcing terms into Eqs. 2 and 3 and quickly evaluated via multiscale expansion pound2022black ; Mathews:2025nyb or near-identity transform algorithms Lynch:2021ogr ; Lynch:2023gpu .
@@
We generate inspiral trajectories by numerically integrating Eqs. 2 and 3 from a set of initial conditions $\{p_{0},e_{0},\Phi_{\phi 0},\Phi_{r0}\}$ with an adaptive ODE solver, returning a sparse trajectory consisting of $\sim 100$ points. Formally, the adiabatic approximation breaks down as the trajectory approaches the separatrix $p_{\mathrm{sep}}(a,e,x_{I})$, which represents the innermost stable orbit in Kerr spacetime stein2020location . As $p\rightarrow p_{\mathrm{sep}}(a,e,x_{I})$, the CO “transitions” across the boundary of the separatrix and plunges into the MBH on timescales where the adiabatic approximation of the inspiral trajectory is no longer valid ori2000transition . In practice, it is sensible to terminate the trajectory at an intermediate point $p_{\mathrm{sep}}+\Delta p_{\mathrm{buf}}$ prior to crossing the separatrix. Not only is this regime better modelled by a multiscale transition-to-plunge expansion (e.g., Refs. Compere:2021iwh ; Compere:2021zfj ; Kuchler:2024esj ; Becker:2024xdi ; Lhost:2024jmw ; Kuchler:2025hwx ), but in our parameterization the forcing functions $\hat{f}^{(0)}_{\alpha}$ diverge as $p\to p_{\rm sep}$ kennefick1998stability ; Hughes:2024tja  [^6] This divergence is due to the Jacobian between the parameters $(p,e,x_{I})$ and the integrals of motion $(E,L_{z},Q)$; see Eq. 7. As discussed in Ref. Hughes:2024tja , one can avoid this issue by changing the parameterization, though at the cost introducing other numerical complications. which leads to numerical instabilities in ODE integration techniques such as Runge–Kutta methods. The point at which one should switch from an inspiral to a transitional model (in terms of matching at first post-adiabatic order) is not yet fully understood for generic trajectories. We can get some guidance by considering quasi-circular orbits, for which it is well understood that the transition away from adiabatic inspiral happens when $p-p_{\rm sep}\approx K(\nu/10^{-5})^{2/5}$ ori2000transition ; Apte:2019txp ; Kuchler:2024esj ; the numerical coefficient $K$ is roughly $0.05$ for Schwarzschild, and varies with primary spin and orbital inclination by a factor of order several tens of percent. As our inspiral model extends to $p-p_{\mathrm{sep}}=2\times 10^{-3}$ (see Section III.1), we expect it will be capable of incorporating such a transitional component in the future, though extending these transition models to eccentric orbits is still an area of active research Becker:2024xdi ; Lhost:2024jmw .
@@
  • For $a\neq 0$, the separatrix no longer takes the simple form $p_{\mathrm{sep}}=6+2e$ and a more costly root-finding operation is often required stein2020location ; Ng:2025maa 777 In principle, it is indeed possible to express the Kerr equatorial separatrix in a closed-form analytical expression $p_{\mathrm{sep}}(a,\,e)$ through the solution of degree 4 polynomials for $p$ (cf. Ref. Compere:2021bkk ). However, the expression is relatively lengthy (involving nested radicals) and it is found that the computational efficiency is similar to that of the pre-existing numerical root finding method. For that reason, we shall not employ this analytical expression in this work..
@@
A description of how the few inspiral generation framework must be modified to address these issues is presented in Section III.1.
@@ II.2 Mode amplitudes for Kerr inspirals
where ${}_{-2}S_{\ell mn}$ are spheroidal harmonics with spin-weight $-2$, and
@@
In analogy to Eq. 4, $\hat{\omega}_{mn}=m\hat{\Omega}_{\phi}+n\hat{\Omega}_{r}$ is the frequency of the mode, and the amplitude $\hat{Z}_{\ell mn}^{\infty}$ is obtained by solving the Teukolsky equation Hughes:2021exa . As the orbital elements $(p,e,x_{I})$ evolve during an inspiral, the dependent quantities $\hat{\omega}_{mn}$, $\hat{Z}_{\ell mn}^{\infty}$ and $S_{\ell mn}$ evolve as well; we indicate this by expressing $A_{\ell mn}$ and $H_{\ell mn}$ as functions of time.
The evolution of the spheroidal harmonics has the potential to greatly increase the cost of computing each mode, since it appears that one must continually recompute these functions as $\hat{\omega}_{mn}$ evolves along the inspiral. To avoid this cost, we use the fact that the spin-weighted spheroidal harmonics ${}_{-2}S_{\ell mn}(\theta;\hat{\omega}_{mn})e^{im\phi}$ can be expanded into spin-weighted spherical harmonics ${}_{-2}Y_{\ell m}(\theta,\phi)$ very efficiently (indeed, the spheroidal harmonics exactly reduce to spherical harmonics when $a=0$ Teukolsky1973 ). These functions do not vary during an inspiral and are inexpensive to compute. We take advantage of this behavior by projecting the amplitudes $A_{\ell mn}$ onto the spherical harmonic basis, allowing us to build each wave mode using angular functions that do not vary along the trajectory. The mode spectrum which results is generally broader in the deep strong field accessible to prograde inspirals (i.e., higher multipoles contribute more significantly to the overall waveform), which places more stringent requirements on the framework used to generate these amplitudes. The remapping procedure and subsequent modifications to the few amplitude module that are required in order to accommodate this increased complexity are described in Section III.2.
@@ III.1 Inspiral trajectory
The accurate and efficient integration of Eqs. 2 and 3 is an essential component of the few framework. However, as pointed out in Section II.1, incorporating the effects of spin presents additional challenges that must be addressed if accuracy and efficiency are to be retained. In this section, we will demonstrate how the existing few framework for trajectory generation (which previously modelled zero-spin systems) can be modified for use as part of a Kerr equatorial eccentric waveform model. We will also discuss general modifications to the trajectory module of few we have made that improve the accuracy and efficiency of trajectories and waveforms, but are not limited in scope to the introduction of MBH spin into the model.
@@ III.1.1 Computation of trajectory fluxes
To evolve the trajectories, we must compute the forcing functions $\hat{f}^{(0)}_{p,e}$ on the righthand side of Eq. 2. At adiabatic order, $\hat{f}^{(0)}_{p,e}$ are related to the flux of energy and angular momentum lost by the system (both to null infinity and through the event horizon of the primary) due to GW emission,
@@
where $\alpha=(p,e)$ and the Jacobian elements ${\partial\alpha}/{\partial E}$ and ${\partial\alpha}/{\partial L}$ are analytic functions of $(a,p,e,x_{I})$ as given in Appendix B of Ref. Hughes:2021exa . The time-averaged GW fluxes $\langle{dE}/{dt}\rangle_{\mathrm{GW}}$ and $\langle{dL}/{dt}\rangle_{\mathrm{GW}}$ are calculated from the Teukolsky equation using techniques in black hole perturbation theory, as described in, e.g., Refs. Fujita:2020zxe ; Skoupy:2021asz ; Hughes:2021exa . We omitted the Carter flux $\langle{dQ}/{dt}\rangle_{\mathrm{GW}}$ in Eq. 7 as it is zero for equatorial orbits. The fluxes are also functions of $(a,p,e,x_{I})$, but must be determined numerically.
In this work, we use the Python library pybhpt Nasipak:2023kuf ; zachary_nasipak_2025_15627818 to compute $\left\langle{dE}/{d{t}}\right\rangle_{\mathrm{GW}}$ and $\left\langle{dL}/{d{t}}\right\rangle_{\mathrm{GW}}$. Fluxes are computed to a requested precision of $10^{-8}$. We chose this value as a balance between computational cost and model accuracy, since multiscale analyses suggest that errors in the fluxes induce $O(\nu^{-1})$ errors in the GW phasing Hinderer:2008dm . The computational cost of each flux calculation ranges from $\sim 1$ to $\sim 1000$ seconds and is highly dependent on the values of $(a,p,e,x_{I})$. As pybhpt performs computations in the frequency-domain, we find that it is most efficient and accurate for orbits where $p>p_{\mathrm{sep}}+1$ and $e<0.1$, since these orbits have a very narrow frequency spectrum. The computational cost of the flux calculations then grows as $p\rightarrow p_{\mathrm{sep}}$ and $e\rightarrow 1$. (See Section B.4 for further details on the timing of flux calculations.) Compared to the rest of the waveform generation, the fluxes are incredibly expensive to compute. In the following section, we describe how this computational cost is circumvented in the few framework.

### v2

@@ II Waveforms for eccentric equatorial inspirals
where $\mu=m_{1}m_{2}/M$ is the reduced mass of the binary, $M=m_{1}+m_{2}$ is its total mass, $d_{\mathrm{L}}$ is the luminosity distance to the source, and $(\theta,\phi)$ are, respectively the source-frame polar and azimuthal viewing angles of the observer. Our notation for the masses in our waveform model deviates from the majority of previous adiabatic waveform models in the GSF literature, e.g., Fujita:2020zxe ; Hughes:2021exa ; Isoyama:2021jjd ; Katz:2021yft ; Nasipak:2023kuf ; Khalvati:2024tzz , which define $\mu=m_{2}$ and $M=m_{1}$. Instead, our mass convention is chosen to align with what is commonly used when modelling comparable-mass binaries, e.g., Blanchet:2013haa ; Blanchet:2023bwj ; Pompili:2023tna ; Albanesi:2025txj . Further discussion of this change in convention is provided in Appendix˜A.
In the following section, we review how Eq.˜1 is constructed in the few framework, highlighting where assumptions made for non-spinning systems must be modified in the extension of this framework to spinning systems.
@@ II.1 Inspiral trajectories into spinning black holes
In order to compute the amplitude and phase of each mode in Eq.˜1, we must first obtain the orbital trajectory of the CO by integrating its equations of motion. Operating in the multiscale framework at adiabatic order Hinderer:2008dm ; Hughes:2016xwf ; pound2022black ; Isoyama:2021jjd , we parameterize the inspiral trajectory at each moment in time by three quasi-Keplerian orbital elements — $p(t)$ (semi-latus rectum), $e(t)$ (eccentricity) and $x_{I}(t)$ (cosine of the orbital inclination)— along with the fundamental orbital phases $\Phi_{A}(t)$ with $A=\{r,\theta,\phi\}$ (up to initial conditions). In this work, we specifically choose these elements such that they satisfy the geodesic relations between semi-latus rectum, eccentricity, inclination and the (dimensionless) orbital frequencies $\hat{\Omega}_{A}(a,p,e,x_{I})$, as described in Lynch:2024hco (built on works Glampedakis:2002ya ; Schmidt:2002qk ; Fujita:2009bp ; KerrGeodesicsPackage ), so that it adheres to the convention set out in the LISA convention document lisa_rosetta_stone_2025 . [^4] Some works treat $p$ as dimensionless (e.g., Fujita:2009bp ), while others take $p$ to scale with $m_{1}$ (e.g., Lynch:2024hco ). In this work, we make all quantities, including $p$, dimensionless and explicitly introduce the dependence on mass through the equations of motion. We discuss our choice of mass convention in Appendix A. The equations of motion then describe how these quantities evolve with respect to time. At adiabatic order, the evolution of the orbital elements is driven solely by the orbit-averaged back-reaction due to GW emission, and the evolution of the fundamental phases is equivalent to that of the tangent geodesic (which changes with respect to time due to the slowly evolving orbital elements). The adiabatic equations of motion are therefore a system of six ordinary differential equations (ODEs) that must be solved numerically. Three of these ODEs are associated with the evolution of the (dimensionless) orbital elements $\alpha\in\{p,e,x_{I}\}$:
@@
These phases can then be related to the $\{r,\theta,\phi\}$ Boyer-Lindquist coordinates of the secondary using the relations found in Ref. Lynch:2024hco . The phase evolution of each mode in Eq.˜1 is then
@@
However, as only modes with $k=0$ contribute to the waveform for equatorial inspirals, we may write $\Phi_{mn}=m\Phi_{\phi}+n\Phi_{r}$ and ignore $\Phi_{\theta}$. By combining Eqs.˜2 and 3, and integrating over a radiation reaction timescale ($\sim M/\nu$), we can see that the truncation in the evolution of the orbital elements leads to an error in the orbital phase of $O(\nu^{0})$ Hinderer:2008dm  [^5] In this work, we are ignoring the GSF effects of transient resonances Flanagan:2010cd , which first enter the inspiral evolution at $\mathcal{O}(\nu^{1/2})$ past the adiabatic order vandeMeent:2013sza ; Berry:2016bit ; Lynch:2024ohd . Transient resonances for equatorial eccentric inspirals are even weaker, entering at $\mathcal{O}(\nu^{1})$ (i.e. first post-adiabatic order) vandeMeent:2014raa and can therefore be neglected in our adiabatic model.. This is sufficient to demonstrate the few framework for IMRIs and EMRIs, and to obtain the quantitive results of the studies performed in this work. Once post-adiabatic corrections to the evolution of the orbital elements are known, they can easily be incorporated as additional forcing terms into Eqs.˜2 and 3 and quickly evaluated via multiscale expansion pound2022black ; Mathews:2025nyb or near-identity transform algorithms Lynch:2021ogr ; Lynch:2023gpu .
@@
We generate inspiral trajectories by numerically integrating Eqs.˜2 and 3 from a set of initial conditions $\{p_{0},e_{0},\Phi_{\phi 0},\Phi_{r0}\}$ with an adaptive ODE solver, returning a sparse trajectory consisting of $\sim 100$ points. Formally, the adiabatic approximation breaks down as the trajectory approaches the separatrix $p_{\mathrm{sep}}(a,e,x_{I})$, which represents the innermost stable orbit in Kerr spacetime stein2020location . As $p\rightarrow p_{\mathrm{sep}}(a,e,x_{I})$, the CO “transitions” across the boundary of the separatrix and plunges into the MBH on timescales where the adiabatic approximation of the inspiral trajectory is no longer valid ori2000transition . In practice, it is sensible to terminate the trajectory at an intermediate point $p_{\mathrm{sep}}+\Delta p_{\mathrm{buf}}$ prior to crossing the separatrix. Not only is this regime better modelled by a multiscale transition-to-plunge expansion (e.g., Refs. Compere:2021iwh ; Compere:2021zfj ; Kuchler:2024esj ; Becker:2024xdi ; Lhost:2024jmw ; Kuchler:2025hwx ), but in our parameterization the forcing functions $\hat{f}^{(0)}_{\alpha}$ diverge as $p\to p_{\rm sep}$ kennefick1998stability ; Hughes:2024tja  [^6] This divergence is due to the Jacobian between the parameters $(p,e,x_{I})$ and the integrals of motion $(E,L_{z},Q)$; see Eq. 7. As discussed in Ref. Hughes:2024tja , one can avoid this issue by changing the parameterization, though at the cost introducing other numerical complications. which leads to numerical instabilities in ODE integration techniques such as Runge–Kutta methods. The point at which one should switch from an inspiral to a transitional model (in terms of matching at first post-adiabatic order) is not yet fully understood for generic trajectories. We can get some guidance by considering quasi-circular orbits, for which it is well understood that the transition away from adiabatic inspiral happens when $p-p_{\rm sep}\approx K(\nu/10^{-5})^{2/5}$ ori2000transition ; Apte:2019txp ; Kuchler:2024esj ; the numerical coefficient $K$ is roughly $0.05$ for Schwarzschild, and varies with primary spin and orbital inclination by a factor of order several tens of percent. As our inspiral model extends to $p-p_{\mathrm{sep}}=2\times 10^{-3}$ (see Section˜III.1), we expect it will be capable of incorporating such a transitional component in the future, though extending these transition models to eccentric orbits is still an area of active research Becker:2024xdi ; Lhost:2024jmw .
@@
  • For $a\neq 0$, the separatrix no longer takes the simple form $p_{\mathrm{sep}}=6+2e$ and a more costly root-finding operation is often required stein2020location ; Ng:2025maa  [^7] In principle, it is indeed possible to express the Kerr equatorial separatrix in a closed-form analytical expression $p_{\mathrm{sep}}(a,\,e)$ through the solution of degree 4 polynomials for $p$ (cf. Ref. Compere:2021bkk ). However, the expression is relatively lengthy (involving nested radicals) and it is found that the computational efficiency is similar to that of the pre-existing numerical root finding method. For that reason, we shall not employ this analytical expression in this work..
@@
A description of how the few inspiral generation framework must be modified to address these issues is presented in Section˜III.1.
@@ II.2 Mode amplitudes for Kerr inspirals
where $_{-2}S_{\ell mn}$ are spheroidal harmonics with spin-weight 2-2, and
@@
In analogy to Eq.˜4, $\hat{\omega}_{mn}=m\hat{\Omega}_{\phi}+n\hat{\Omega}_{r}$ is the frequency of the mode, and the amplitude $\hat{Z}_{\ell mn}^{\infty}$ is obtained by solving the Teukolsky equation Hughes:2021exa . As the orbital elements $(p,e,x_{I})$ evolve during an inspiral, the dependent quantities $\hat{\omega}_{mn}$, $\hat{Z}_{\ell mn}^{\infty}$ and $S_{\ell mn}$ evolve as well; we indicate this by expressing $A_{\ell mn}$ and $H_{\ell mn}$ as functions of time.
The evolution of the spheroidal harmonics has the potential to greatly increase the cost of computing each mode, since it appears that one must continually recompute these functions as $\hat{\omega}_{mn}$ evolves along the inspiral. To avoid this cost, we use the fact that the spin-weighted spheroidal harmonics $_{-2}S_{\ell mn}(\theta;\hat{\omega}_{mn})e^{im\phi}$ can be expanded into spin-weighted spherical harmonics $_{-2}Y_{\ell m}(\theta,\phi)$ very efficiently (indeed, the spheroidal harmonics exactly reduce to spherical harmonics when $a=0$ Teukolsky1973 ). These functions do not vary during an inspiral and are inexpensive to compute. We take advantage of this behavior by projecting the amplitudes $A_{\ell mn}$ onto the spherical harmonic basis, allowing us to build each wave mode using angular functions that do not vary along the trajectory. The mode spectrum which results is generally broader in the deep strong field accessible to prograde inspirals (i.e., higher multipoles contribute more significantly to the overall waveform), which places more stringent requirements on the framework used to generate these amplitudes. The remapping procedure and subsequent modifications to the few amplitude module that are required in order to accommodate this increased complexity are described in Section˜III.2.
@@ III.1 Inspiral trajectory
The accurate and efficient integration of Eqs.˜2 and 3 is an essential component of the few framework. However, as pointed out in Section˜II.1, incorporating the effects of spin presents additional challenges that must be addressed if accuracy and efficiency are to be retained. In this section, we will demonstrate how the existing few framework for trajectory generation (which previously modelled zero-spin systems) can be modified for use as part of a Kerr equatorial eccentric waveform model. We will also discuss general modifications to the trajectory module of few we have made that improve the accuracy and efficiency of trajectories and waveforms, but are not limited in scope to the introduction of MBH spin into the model.
@@ III.1.1 Computation of trajectory fluxes
To evolve the trajectories, we must compute the forcing functions $\hat{f}^{(0)}_{p,e}$ on the righthand side of Eq.˜2. At adiabatic order, $\hat{f}^{(0)}_{p,e}$ are related to the flux of energy and angular momentum lost by the system (both to null infinity and through the event horizon of the primary) due to GW emission,
@@
where $\alpha=(p,e)$ and the Jacobian elements ${\partial\alpha}/{\partial E}$ and ${\partial\alpha}/{\partial L}$ are analytic functions of $(a,p,e,x_{I})$ as given in Appendix B of Ref. Hughes:2021exa . The time-averaged GW fluxes $\langle{dE}/{dt}\rangle_{\mathrm{GW}}$ and $\langle{dL}/{dt}\rangle_{\mathrm{GW}}$ are calculated from the Teukolsky equation using techniques in black hole perturbation theory, as described in, e.g., Refs. Fujita:2020zxe ; Skoupy:2021asz ; Hughes:2021exa . We omitted the Carter flux $\langle{dQ}/{dt}\rangle_{\mathrm{GW}}$ in Eq.˜7 as it is zero for equatorial orbits. The fluxes are also functions of $(a,p,e,x_{I})$, but must be determined numerically.
In this work, we use the Python library pybhpt Nasipak:2023kuf ; zachary_nasipak_2025_15627818 to compute $\left\langle{dE}/{d{t}}\right\rangle_{\mathrm{GW}}$ and $\left\langle{dL}/{d{t}}\right\rangle_{\mathrm{GW}}$. Fluxes are computed to a requested precision of $10^{-8}$. We chose this value as a balance between computational cost and model accuracy, since multiscale analyses suggest that errors in the fluxes induce $O(\nu^{-1})$ errors in the GW phasing Hinderer:2008dm . The computational cost of each flux calculation ranges from $\sim 1$ to $\sim 1000$ seconds and is highly dependent on the values of $(a,p,e,x_{I})$. As pybhpt performs computations in the frequency-domain, we find that it is most efficient and accurate for orbits where $p>p_{\mathrm{sep}}+1$ and $e<0.1$, since these orbits have a very narrow frequency spectrum. The computational cost of the flux calculations then grows as $p\rightarrow p_{\mathrm{sep}}$ and $e\rightarrow 1$. (See Section˜B.4 for further details on the timing of flux calculations.) Compared to the rest of the waveform generation, the fluxes are incredibly expensive to compute. In the following section, we describe how this computational cost is circumvented in the few framework.